OpenStudy (anonymous):

The circumfrence of a circle is 18(pie symbol) find the radius and area of the circle

OpenStudy (anonymous):

\[18\pi= \pi d\] d= diameter. Solve for d.

OpenStudy (anonymous):

\[2\pi r=\pi d\]

OpenStudy (anonymous):

Once oyu found , you can divide by 2.

OpenStudy (anonymous):

The formula for the circumference of a circle is 2r(pi). You are given that the circumference is 18(pi). Radius is 9 units, so find area. Area = \[(\pi)(r ^{2})\]

OpenStudy (anonymous):

Once you*

OpenStudy (anonymous):

found d*

OpenStudy (anonymous):

@hsrikako so the area is 3.14 * 18

OpenStudy (anonymous):

No, your radius is 9. So Area is 3.14 * (9^2) or 3.14*(81)